# Affine Chord Sobolev Inequalities from Integral Geometry

### v1

@@ 9.2. Results for $s$-concave functions
$\lim_{\alpha\rightarrow-1^{+}}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u)=(n+\frac{1}{s})\|f\|_{1}\rho_{\Pi^{*}f}(u).$
which yields the right inclusion. ∎
When $s=\infty$, Theorem 9.8 reduces to the geometric case in Theorem 9.1 by Gardner and Zhang. In fact, following the argument of our proof, for a log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, if one could show that $g(tu)^{1/n}$ is concave, then the following chain of inclusions would hold:
${\rm R}_{\infty}f\subset c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f\subset n\|f\|_{1}\Pi^{*}f,$
for $1<\alpha<\beta<\infty$.
This would yield precisely the functional counterpart of Theorem 9.1 that one would expect. However, $g(tu)^{1/n}$ is not necessarily concave for a general log-concave function $f$. A counterexample will be provided below.
Let $n=1$ and
[…]
It is clear that $\log f(x)=\log x+\log 2$ for $t\in[0,1]$, which is concave. Then we have
[…]
Note that
$\frac{1}{4}=g(\frac{1}{2})<\frac{1}{2}(g(0)+g(1))=\frac{1}{2},$
which shows that $g$ is not concave. In this case, we consider $\beta=1$ and $\alpha=1/2$ and obtain that
$c_{1,1}{\rm R}_{1}f=\Big[-\frac{2}{3},\frac{2}{3}\Big],\quad c_{1,\frac{1}{2}}{\rm R}_{\frac{1}{2}}f=\Big[-\frac{16}{25},\frac{16}{{25}}\Big].$
The following theorem describes the sharpness of the inclusion result in Theorem 5.
@@
Equality holds in (9.7) if $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$ where $x_{0}\in\mathbb{R}^{n}$, $c>0$ and $\triangle$ is a simplex containing the origin.
@@
Note that ${\rm R}_{\alpha}f$ is invariant under scaling and translation of $f$. Without loss of generality, we may assume $f(x)=(1-\|x\|_{\triangle})_{+}^{1/s}$. A direct computation then gives
$\|f\|_{1}=B(\frac{1}{s},n+1)\frac{|\triangle|}{s}.$
We use the same notation for $g(tu)$ with a fixed $u\in\mathbb{S}^{n-1}$, and we have
$g(tu)$ $=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
$=\int_{0}^{1}|\{f\geq r\}\cap(\{f\geq r\}+tu)|dr$
$=\int_{0}^{1}\frac{1}{s}r^{\frac{1}{s}-1}|(1-r)\triangle\cap((1-r)\triangle+tu)|dr$
$=\int_{0}^{1}\frac{1}{s}r^{\frac{1}{s}-1}(1-r)^{n}|\triangle\cap(\triangle+\frac{t}{1-r}u)|dr.$
By [18, Lemma 3.4], $|\triangle\cap(\triangle+tu)|^{1/n}$ is linear in $t$. Therefore, there exists $\lambda>0$ such that
$|\triangle\cap(\triangle+\frac{t}{1-r})u|=|\triangle|(1-\lambda\frac{t}{1-r})_{+}^{n}.$
A straightforward, though somewhat tedious, calculation then shows that
$g(tu)=\frac{1}{s}B(\frac{1}{s},n+1)|\triangle|(1-\lambda t)^{\frac{ns+1}{s}}.$
We set $s^{\prime}=s/(ns+1)$. Then, for $1<\alpha<0$ and $\alpha>0$, by taking $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$ in Lemma 9.2, we obtain
$c_{n,\alpha}(s){\rm R}_{\alpha}f=\frac{1}{s^{\prime}}\zeta(\alpha)=\frac{1}{s^{\prime}}\Big(\frac{\int_{0}^{\infty}t^{\alpha-1}(1-\lambda t)_{+}^{1/s^{\prime}}dt}{\int_{0}^{\infty}t^{\alpha-1}(1-s^{\prime}t)^{1/s^{\prime}}dt}\Big)^{1/\alpha}=\frac{1}{\lambda},$
where the first equality is from thex calculation in Theorem 9.8.
@@
Letting $\alpha\rightarrow-1$ and $\beta\rightarrow\infty$, we have
@@
We conclude the section by showing the necessary condition for the equality cases in Theorem 5. The following lemma is the key result for the equality characterization.
@@
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ non-zero and $s$-concave with compact support. If
$g(tu)^{\frac{s}{ns+1}}=\Big(\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx\Big)^{\frac{s}{ns+1}}$
is linear in $t$ for any fixed $u\in\mathbb{S}^{n-1}$, then $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$, where $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
@@
We assume, without loss of generality, that $\|f\|_{\infty}=1$. Since
$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x-tu)\}dx,$
for convenience of notations in the following proof, we set $N_{f}(x,t)=\min\{f(x),f(x-tu)\}$ here. The functions $F,G,H$ are as defined by
$F(x)=N_{f}(x,0),~~G(x)=N_{f}(x,t),~~H(x)=N_{f}(x,\lambda t),$
which satisfies that $H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y))$.
The linearity of $g(tu)^{\frac{s}{ns+1}}$ then implies that
@@
That is, we have the equality in the Borell-Brascamp-Lieb inequality for $F,G,H$.
By Theorem 2.4, there exists $c_{0}=\Big(\frac{|{\rm supp}(G)|}{|{\rm supp}(F)|}\Big)^{1/n}$ and $y_{0}\in\mathbb{R}^{n}$ such that
${\rm supp}(G)=c_{0}\,{\rm supp}(F)+y_{0},$ (9.9)
while the definition of $G$ shows
${\rm supp}(G)={\rm supp}(F)\cap({\rm supp}(F)+tu).$ (9.10)
Since $u\in\mathbb{S}^{n-1}$ is arbitrary fixed, (9.9) and (9.10) shows that ${\rm supp}(F)\cap({\rm supp}(F)+tu)$ is homothetic to ${\rm supp}(F)$ for all $u\in\mathbb{S}^{n-1}$. By [18, Lemma 3.4], we have ${\rm supp}(F)$ is a simplex $\triangle$, where we assume that $|\triangle|$ is the same as the volume of the standard simplex $\triangle_{n}$ for convenience.
In fact, we only need to consider when $\triangle$ is the standard simplex, that is,
@@
Otherwise, there exists $\phi\in{\rm SL}(n)$ such that ${\rm supp}(f\circ\phi^{-1})=\triangle_{n}$ and we consider $f\circ\phi^{-1}$ instead of $f$. In this case, by (9.9) and (9.10), we have
$c_{0}(t)=1-t\sum_{i=1}^{n}u_{i},~~\text{and}~~y_{0}(t)=t(u_{1}^{+},\ldots,u_{n}^{+}),$
where $u_{i}^{+}=\max\{0,u_{i}\}$ and we denote by $a_{0}=\sum_{i=1}^{n}u_{i}$. Therefore, we can choose $u\in\mathbb{S}^{n-1}$ with $u_{i}\geq 0$, such that $y_{0}(t)=tu$.
Moreover, by Theorem 2.4, there exists a $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(F)\rightarrow[0,\infty)$ with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$ satisfying,
$\left\{\begin{aligned} &F(x)=\Phi(x)&&,\\ &G(c_{0}x+tu)=c_{0}^{\frac{1}{s}}\Phi(x)&&,\\ &H((1-\lambda+\lambda c_{0})x+\lambda tu)=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{0}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&.\end{aligned}\right$.
In fact, $\Phi(x)=F(x)=f(x)$ by the definition of $F$. Then we have
$G((1-a_{0}t)x+tu)^{s}=\min\{f((1-a_{0}t)x)^{s},f((1-a_{0}t)x+tu)^{s}\}=(1-a_{0}t)f(x)^{s}.$
Note that $(1-a_{0}t)f(x)^{s}\neq f((1-a_{0}t)x)^{s}$ since $f^{s}$ is concave. Therefore,
$f((1-a_{o}t)x+tu)^{s}=(1-a_{0}t)f(x)^{s}$ (9.11)
holds for $0\leq t\leq 1/\sum u_{i}$. Equation (9.11), together with $f\geq 0$ and the $s$-concavity of $f$ , imply that
$f\Big(\frac{u}{a_{0}}\Big)^{s}=0.$ (9.12)
We set $\bar{x}=(\bar{x}_{1},\ldots,\bar{x}_{n})\in\triangle_{n}$ such that $f(\bar{x})=\|f\|_{\infty}=1$. Note that if $\sum_{i}\bar{x}_{i}\in(0,1)$, we will have
$f\Big(\frac{\bar{x}}{\sum_{i}\bar{x}_{i}}\Big)^{s}=0\geq(1-\sum_{i}\bar{x}_{i})f(\bar{x})^{s}+\Big(\sum_{i}\bar{x}_{i}\Big)f(o)^{s}\geq 1-\sum_{i}\bar{x}_{i}>0,$ (9.13)
which cause a contradiction. Thus we have $\sum_{i}\bar{x}_{i}=1$ or $\bar{x}=o$. If $\sum_{i}\bar{x}_{i}=1$, equation (9.12) implies that $f(\bar{x})=0$ which contradicts to the choice of $\bar{x}$. Thus we have $\bar{x}=o$. Letting $x=\bar{x}=o$ in (9.11) then gives
$f(tu)^{s}=1-a_{0}t=1-\|tu\|_{\triangle_{n}}.$
Since ${\rm supp}(f)=\triangle_{n}$, we then have
$f(x)=(1-\|x\|_{\triangle_{n}})_{+}^{1/s},$
which concludes the proof. ∎
We conclude this section by the following theorem, which provides the full equality characterization in Theorem 5.
@@
Let $s>0$ and $c_{n,\alpha}(s)$ be defined in Theorem 5. Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, and $s$-concave function with compact support. For $1<\alpha<\beta<\infty$, if
@@
then $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$ where $c>0$, $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.

### v2

@@ 6.2. Results for $s$-concave functions
$\lim_{\alpha\rightarrow-1^{+}}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u)=(n+\frac{1}{s})\|f\|_{1}\rho_{\Pi^{*}f}(u),$
which yields the right-most inclusion. ∎
The following theorem demonstrates the sharpness of (6.6).
@@
Let $s>0$. If $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $a=\|f\|_{\infty}>0$ and $\triangle\subset\mathbb{R}^{n}$ is a simplex containing $x_{0}$, then
where $1<\alpha<\beta<\infty$ and $c_{n,\alpha}(s)$ is given in Theorem 6.8.
@@
By Lemma 5.3, ${\rm R}_{\alpha}f$ is invariant under scaling and translation. We may assume, without loss of generality, that $x_{0}=o$ and $a=1$. Then
$\|f\|_{1}=\int_{0}^{1}|\{f\geq t\}|dt=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}|\{f^{s}\geq t\}|dt,$
where $\{f^{s}\geq t\}=(1-t)\triangle$. Hence
$\|f\|_{1}=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}(1-t)^{n}|\triangle|dt=B(\frac{1}{s},n+1)\frac{|\triangle|}{s}.$
Moreover, we obtain
$g(ru)$ $=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$
$=\int_{0}^{1}|\{f\geq t\}\cap(\{f\geq t\}+ru)|dt$
$=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}|(1-t)\triangle\cap((1-t)\triangle+ru)|dt$
$=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}(1-t)^{n}\Big|\triangle\cap(\triangle+\frac{r}{1-t}u)\Big|dt.$
By [18, Lemma 3.4], $|\triangle\cap(\triangle+ru)|^{1/n}$ is an affine function with respect to $r>0$. Therefore, there exists $\lambda>0$ such that
$\Big|\triangle\cap(\triangle+\frac{r}{1-t})u\Big|=|\triangle|\Big(1-\lambda\frac{r}{1-t}\Big)_{+}^{n},$
where $\lambda$ depends on $u$.
A straightforward calculation shows that
$g(ru)=\frac{1}{s}B(\frac{1}{s},n+1)|\triangle|(1-\lambda r)_{+}^{\frac{ns+1}{s}}.$
We set $s^{\prime}=s/(ns+1)$. For $1<\alpha<0$ and $\alpha>0$, by taking $\omega(r)=g(o)(1-s^{\prime}r)_{+}^{1/s^{\prime}}$ and $\varphi(r)=\frac{1}{s^{\prime}}(1-(g(ru)/g(o))^{s^{\prime}})$ in Lemma 6.2, we obtain
$c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u)=\frac{1}{s^{\prime}}\zeta(\alpha)=\frac{1}{s^{\prime}}\Big(\frac{\int_{0}^{\infty}r^{\alpha-1}(1-\lambda r)_{+}^{1/s^{\prime}}dr}{\int_{0}^{\infty}r^{\alpha-1}(1-s^{\prime}r)_{+}^{1/s^{\prime}}dr}\Big)^{1/\alpha}=\frac{1}{\lambda},$
where the first equality is from the calculation in the proof of Theorem 6.8.
@@
Letting $\alpha\rightarrow-1$ and $\beta\rightarrow\infty$ completes the proof. ∎
To establish the equality characterization, we first need an auxiliary lemma based on the equality case of the Borell–Brascamp–Lieb inequality (2.4). Before stating this result, we collect some preliminary facts that will be used later in the argument. Throughout, ${\rm cl}K$ denotes the closure of a set $K$.
Lemma 6.10.
Let $f:\mathbb{R}\rightarrow[0,\infty)$ be a concave function on $(0,c)$ with ${\rm supp}(f)=[0,c]$, and define
$K={\rm cl}\{(x,y)\in\mathbb{R}^{2}:x\in(0,c)~\text{and}~0\leq y\leq f(x)\}.$
Let $e_{1}=(1,0)\in\mathbb{R}^{2}$. For all $r\in(0,c)$, if either
$K\cap(K-re_{1})=(1-\tfrac{r}{c})K\quad\text{or}\quad K\cap(K+re_{1})=(1-\tfrac{r}{c})K+re_{1},$
then $K$ is a triangle.
Proof.
We assume that $K\cap(K-re_{1})=(1-\frac{r}{c})K$ and prove the statement, the other case is similar. Let
$x_{0}=\inf\{x\in[0,c]:f(x)=\|f\|_{\infty}\}~\text{and}~x_{1}=\sup\{x\in[0,c]:f(x)=\|f\|_{\infty}\}.$
We first aim to show that $x_{0}=x_{1}$. If $x_{1}>x_{0}$, we choose $r>0$ to be sufficiently small such that $K\cap(K-re_{1})\neq\emptyset$ and $r<x_{1}-x_{0}$. Then the identity $K\cap(K-re_{1})=(1-\frac{r}{c})K$ implies that
$x_{1}-x_{0}-r=(1-\tfrac{r}{c})(x_{1}-x_{0}).$
It follows that either $x_{1}-x_{0}=c$ or $x_{1}-x_{0}=0$. Since the case $x_{1}-x_{0}=c$ would imply that $K$ is a rectangle, contradicting the assumption $K\cap(K-re_{1})=(1-\frac{r}{c})K$, we conclude that $x_{0}=x_{1}$.
Next we show that $f(0)=0$ if $x_{0}\neq 0$. Since $f$ is concave and hence continuous on $(0,c)$, we have $f(0^{-})=\lim_{x\rightarrow 0^{-}}f(x)=0$ and $f(0^{+})=\lim_{x\rightarrow 0^{+}}f(x)\geq f(0)\geq 0$. Thus, it suffices to show that $f(0^{+})=0$. Since $x_{0}$ is the unique maximizer, $f$ is monotone increasing on $(0,x_{0})$, and therefore $f(r)\geq f(0^{+})$ for $r>0$. Then
$\big\{(0,y):y\in[0,f(0^{+})]\big\}\subset K\cap(K-re_{1})=(1-\tfrac{r}{c})K,$
with $r$ sufficiently small. Consequently, $f(0^{+})=(1-\frac{r}{c})f(0^{+})$, which implies that $f(0)=0$. Similarly, one obtains $f(c)=0$ provided that $x_{0}\neq c$.
We now prove that $K$ is a triangle provided that $x_{0}\in(0,c)$. In this case, we have $f(0)=f(c)=0$. Since $x_{0}$ is the unique maximizer, $f$ is monotone increasing on $(0,x_{0})$ and monotone decreasing on $(x_{0},c)$. The identity $K\cap(K-re_{1})=(1-\frac{r}{c})K$ then implies that
$f\big((1-\tfrac{r}{c})x_{0}\big)=(1-\tfrac{r}{c})f(x_{0}),$
and
$f\big((1-\tfrac{r}{c})x_{0}+r\big)=(1-\tfrac{r}{c})f(x_{0}).$
This shows that $f$ is linear on $0,x_{0}]$ and affine on $x_{0},c]$, which implies that $K$ is a triangle. For $x_{0}=0$ or $x_{0}=c$, the proof is similar. ∎
Next we aim to extend Lemma 6.10 to concave functions on $\mathbb{R}^{n}$. For a non-negative, concave function $f\in L^{1}(\mathbb{R}^{n})$, we denote by
${\rm H}(f)={\rm cl}\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}~{\rm supp}(f)~\text{and}~0\leq l\leq f(x)\}$ (6.8)
the hypograph of $f$. Here ${\rm int}K$ denotes the interior of a set $K$. Let $e_{1},\ldots,e_{n}$ be an orthonormal basis of $\mathbb{R}^{n}$. To simplify the notation, we only consider the function whose support is the standard simplex defined by
$\triangle_{n}=\Big\{(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}:x_{i}\geq 0~\text{for}~i=1,\ldots,n,~\text{and}~\sum_{i=1}^{n}x_{n}\leq 1\Big\}.$ (6.9)
Recall that $a^{+}=\max\{a,0\}$. Here we also denote $a^{-}=\max\{-a,0\}$.
For $u=(u_{1},\ldots,u_{n})\in\mathbb{S}^{n-1}$, we denote by $u^{+}=(u_{1}^{+},\ldots,u_{n}^{+})$ and $u^{-}=(u_{1}^{-},\ldots,u_{n}^{-})$. Moreover, we set
$a_{0}=\sum_{i=1}^{n}u_{i}^{+},~b_{0}=\sum_{i=1}^{n}u_{i}^{-},~\text{and}~c_{0}=\max\{a_{0},b_{0}\}.$ (6.10)
One can verify geometrically that
$\triangle_{n}\cap(\triangle_{n}+ru)=(1-c_{0}r)\triangle_{n}+ru^{+},$
for any $u\in\mathbb{S}^{n-1}$.
The next result extends Lemma 6.10 to higher dimensional cases. Since we identify $\mathbb{R}^{n}$ with the hyperplane $\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n+1}$, we sometimes write $x\in\mathbb{R}^{n}$ also for the corresponding point in $\mathbb{R}^{n+1}$, instead of $(x,0)$.
Lemma 6.11.
Let $u\in\mathbb{S}^{n-1}$ be fixed, and $c_{0}$ be defined in (6.10). Suppose $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is concave and ${\rm supp}(f)=\triangle_{n}$, which is the standard simplex given by (6.9). For all $r\in(0,\frac{1}{c_{0}})$ and $u\in\mathbb{S}^{n-1}$, if
${\rm H}(f)\cap\big({\rm H}(f)+ru\big)=(1-c_{0}r){\rm H}(f)+ru^{+},$ (6.11)
then $f(x)=a(1-\|x-x_{0}\|_{\triangle_{n}-x_{0}})_{+}$ with $a=\|f\|_{\infty}$ and $x_{0}\in\triangle_{n}$.
Proof.
We may assume $\|f\|_{\infty}=1$ without loss of generality. We aim to show that $f$ is linear on $o,x_{0}]$ and affine on $x_{0},e_{i}]$ for all $i\in\{1,\ldots,n\}$, where $[x,y]$ denotes the line segment between $x$ and $y$. The case where $x_{0}=o$ or $x_{0}=e_{i}$ is analogous.
We first consider $f$ on $o,x_{0}]$. Let $u=-\frac{x_{0}}{|x_{0}|}$ and thus we have $u^{+}=o$. We choose the $2$-dimensional subspace $T$ spanned by $u$ and $(0,\ldots,0,1)\in\mathbb{R}^{n+1}$. Then (6.11) implies that
$K\cap(K+ru)=(1-c_{0}r)K,$
where $K={\rm H}(f)\cap T$. By Lemma 6.10, we obtain the linearity of $f$ on $o,x_{0}]$.
Next we consider $f$ on $x_{0},e_{i}]$ for any fixed $i\in\{1,\ldots,n\}$. In this case, we let $u=\frac{x_{0}-e_{i}}{|x_{0}-e_{i}|}$ and then
$u^{+}=\frac{1}{|e_{i}-x_{0}|}(x_{0}^{(1)},\ldots,x_{0}^{(i-1)},0,x_{0}^{(i+1)},\ldots,x_{0}^{(n)})~\text{and}~c_{0}=\frac{1-x_{0}^{(i)}}{|x_{0}-e_{i}|}.$
We choose the $2$-dimensional subspace $T$ spanned by $u$ and $(x_{0},1)$. Restricting (6.11) to $T$, we obtain
$\big({\rm H}(f)\cap T\big)\cap\big(({\rm H}(f)\cap T)+ru\big)=(1-c_{0}r){\rm H}(f)\cap T+ru^{+}.$ (6.12)
Set $K={\rm H}(f)\cap T\subset T$. The factor $(1-c_{0}r)$ in (6.12) corresponds to a dilation in $T$, and hence
$K\cap(K+ru)=(1-c_{0}r)\cdot K+ru.$
Here, $(1-c_{0}r)\cdot K$ denotes the dilation of $K$ in $T$ with respect to a chosen point $e_{i}$ in $T$, which serves as the origin for the scaling in the affine subspace $T$. By Lemma 6.10, $f$ is affine on $x_{0},e_{i}]$.
Therefore, we have that ${\rm H}(f)$ is a simplex in $\mathbb{R}^{n+1}$ with vertices $o,e_{1},\ldots,e_{n},(x_{0},1)$ and attains 0 on the boundary point, that is,
$f(x)=(1-\|x-x_{0}\|_{\triangle_{n}-x_{0}})_{+},$
@@
We also need the following basic, but non-trivial result.
@@
Suppose $A,B\subset\mathbb{R}^{n}$ are convex sets with ${\rm int}A\cap{\rm int}B\neq\emptyset$. Then ${\rm cl}(A\cap B)={\rm cl}A\cap{\rm cl}B$.
@@
Since $A\cap B\subset A$ and $A\cap B\subset B$, taking closures yields ${\rm cl}(A\cap B)\subset{\rm cl}A\cap{\rm cl}B$. Let $x\in{\rm cl}A\cap{\rm cl}B$. We choose $x_{0}\in{\rm int}A\cap{\rm int}B$ and let
$x_{\lambda}=(1-\lambda)x_{0}+\lambda x,~\text{for}~\lambda\in[0,1).$
By [42, Lemma 1.1.9], we have $x_{\lambda}\in{\rm int}A\cap{\rm int}B$. Since $x_{\lambda}\rightarrow x$ as $\lambda\rightarrow 1^{-}$, we obtain $x\in{\rm cl}({\rm int}A\cap{\rm int}B)\subset{\rm cl}(A\cap B)$. ∎
The following result associates the equality characterization of the Borell-Brascamp-Lieb inequality with (6.11) in Lemma 6.11.
Lemma 6.13.
Let $u\in\mathbb{S}^{n-1}$ be fixed, and $c_{0}$ be defined in (6.10). Suppose $f$ is a non-negative, non-zero, concave function on $\mathbb{R}^{n}$ with ${\rm supp}(f)=\triangle_{n}$. If
$\min\{f((1-c_{0}r)x+ru^{+}),f((1-c_{0}r)x+ru^{-})\}=(1-c_{0}r)f(x),$ (6.13)
for all $x\in{\rm int}\triangle_{n}$, then
${\rm H}(f)\cap\big({\rm H}(f)+ru\big)=(1-c_{0}r){\rm H}(f)+ru^{+},$
for $r\in(0,\frac{1}{c_{0}})$, where ${\rm H}(f)$ denotes the hypograph of $f$ as defined in (6.8).
Proof.
Let $c_{r}=1-c_{0}r$. For clarity of notation, we denote by
$P=\{(x,l)\in{\mathbb{R}^{n+1}}:x\in{\rm int}\triangle_{n},~0\leq l\leq f(x)\},$
and $m(x)=\min\{f((1-c_{0}r)x+ru^{+}),f((1-c_{0}r)x+ru^{-})\}$. Then ${\rm H}(f)={\rm cl}P$.
We first have
$\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}\triangle_{n},~0\leq l\leq c_{r}f(x)\}=\phi_{r}P,$ (6.14)
where $\phi_{r}={\rm diag}(1,\ldots,1,c_{r})\in{\rm GL}(n+1)$.
To consider the hypograph of $m(x)$, we first denote
$P_{1}$ $=\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}\triangle_{n},~0\leq l\leq f((c_{r}x+ru^{+})\}$ (6.15)
$~=\psi_{r}(\{(x-ru^{+},l):x\in{\rm int}\triangle_{n},~0\leq l\leq f(x)\})$
$~=\psi_{r}(P-ru^{+}),$
where $\psi_{r}={\rm diag}(c_{r}^{-1},\ldots,c_{r}^{-1},1)$. Similarly, we denote
$P_{2}=\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}\triangle_{n},~0\leq l\leq f((c_{r}x+ru^{-})\}=\psi_{r}(P-ru^{-}).$ (6.16)
Note that $\psi_{r}=c_{r}^{-1}\phi_{r}$ and $(A-x)\cap(A-y)=A\cap(A+(x-y))-x$ holds for any set AA and $x,y$. Combining this with (6.14), (6.15), (6.16) and the assumption (6.13), we obtain
$P\cap(P+ru)=c_{r}P+ru^{+}.$ (6.17)
Recall that $r\in(0,\frac{1}{c_{0}})$, which yields that ${\rm int}(P+ru)\cap{\rm int}P\neq\emptyset$. Then Lemma 6.12 implies that
${\rm cl}\big(P\cap(P+ru)\big)={\rm cl}P\cap({\rm cl}P+ru).$
Since ${\rm cl}P={\rm H}(f)$, we conclude the proof by taking closures on both sides of (6.17). ∎
Combining Lemma 6.13 with Lemma 6.11, we obtain the following crucial corollary.
Corollary 6.14.
For $u\in\mathbb{S}^{n-1}$, let $c_{0}$ be defined in (6.10). Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a non-zero, concave function with ${\rm supp}(f)=\triangle_{n}$. For $r\in(0,\frac{1}{c_{0}})$, if
$\min\{f((1-c_{0}r)x),f((1-c_{0}r)x+ru)\}=(1-c_{0}r)f(x),$
for all $x\in{\rm int}\triangle_{n}$ and $u\in\mathbb{S}^{n-1}$, then $f(x)=a(1-\|x-x_{0}\|_{\triangle_{n}-x_{0}})_{+}$ with $a=\|f\|_{\infty}$ and $x_{0}\in\triangle_{n}$.
We are now in a position to prove the following key result for the equality characterization in Theorem 3.
Lemma 6.15.
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, $s$-concave function with compact support. For $u\in\mathbb{S}^{n-1}$, let
$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx.$
If there is a constant $\lambda>0$ such that $g(ru)=\|f\|_{1}(1-\lambda r)_{+}^{\frac{ns+1}{s}}$ for $r>0$, then $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
Proof.
We assume $\|f\|_{\infty}=1$ without loss of generality. Since g(ru)=g(ru)g(ru)=g(-ru), we set
$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x-ru)\}dx,$
and $N_{f}(x,r)=\min\{f(x),f(x-ru)\}$ to simplify notation. Let $r\geq 0$ and $\lambda\in[0,1]$ be fixed, and define the functions $F,G,H:\mathbb{R}^{n}\to\mathbb{R}$ by
$F(x)=N_{f}(x,0)=f(x),~~G(x)=N_{f}(x,r),~~H(x)=N_{f}(x,\lambda r),$
which then satisfy
$H((1-\lambda)x+\lambda y)\geq\Big((1-\lambda)F(x)^{s}+\lambda G(y)^{s}\Big)^{1/s}=\mathcal{M}_{\lambda}^{s}(F(x),G(y))$
for all $x,y\in\mathbb{R}^{n}$ .
Since $r\mapsto g(ru)^{\frac{s}{ns+1}}$ is an affine function with respect to $r$, then
@@
Hence, equality holds in the Borell-Brascamp-Lieb inequality (2.4) for $F,G,H$.
By Theorem 2.4, there exists $c_{r}=\Big(\frac{|{\rm supp}(G)|}{|{\rm supp}(f)|}\Big)^{1/n}$ and $y_{r}\in\mathbb{R}^{n}$ such that
${\rm supp}(G)=c_{r}\,{\rm supp}(f)+y_{r},$ (6.18)
while the definition of $G$ gives
${\rm supp}(G)={\rm supp}(f)\cap({\rm supp}(f)+ru).$ (6.19)
Since $u\in\mathbb{S}^{n-1}$ is arbitrarily fixed, (6.18) and (6.19) imply that ${\rm supp}(f)\cap({\rm supp}(f)+ru)$ is homothetic to ${\rm supp}(f)$ for all $u\in\mathbb{S}^{n-1}$. By [18, Lemma 3.4], ${\rm supp}(f)$ is a simplex denoted by $\triangle$.
By Lemma 5.3, ${\rm R}_{\alpha}f$ is invariant under translations and ${\rm R}_{\alpha}(f\circ\phi^{-1})=\phi{\rm R}_{\alpha}f$ for all $\phi\in{\rm GL}(n)$. It suffices to consider the case where $\triangle$ is the standard simplex, that is,
@@
By (6.18) and (6.19), we have
$c_{r}=1-c_{0}r,~~\text{and}~~y_{r}=ru^{+},$
where $c_{0}$ is defined by (6.10).
Moreover, by Theorem 2.4, there exists a $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(F)\rightarrow[0,\infty)$, with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$, satisfying
$\left\{\begin{aligned} &F(x)=\Phi(x)&&,\\ &G(c_{r}x+y_{r})=c_{r}^{\frac{1}{s}}\Phi(x)&&,\\ &H((1-\lambda+\lambda c_{0})x+\lambda y_{r})=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{r}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&,\end{aligned}\right$.
for almost all $x\in\triangle_{n}$. By the definition of $F$, we have $\Phi(x)=F(x)=f(x)$, and
$G((1-c_{0}r)x+y_{r})^{s}=\min\{f((1-c_{0}r)x+ru^{-})^{s},f((1-c_{0}r)x+ru^{+})^{s}\}=(1-c_{0}r)f(x)^{s}.$ (6.20)
Since $f^{s}$ is concave and hence continuous on ${\rm int}\triangle_{n}$, the equality (6.20) holds for all $x\in{\rm int}\triangle_{n}$. By applying Corollary 6.14, we complete the proof. ∎
As a consequence, Lemma 6.15 derives the following theorem, which provides the full equality characterization in Theorem 3.
@@
Let $s>0$ and $c_{n,\alpha}(s)$ be as defined in Theorem 3. Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. For $1<\alpha<\beta<\infty$, if
@@
then $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.